# Learning to Project for Cross-Task Knowledge Distillation

### v1

@@ 3.3 Setup and Training loss
Our cross-task knowledge distillation pipeline is shown in figure 0(b). It consists of a trainable student model, which is to be trained on a given target task, and a frozen teacher model, which is pre-trained on a different task. This setup is in contrast to the traditional same-task knowledge distillation setting, which is shown in figure 0(a). Both the student and teacher models receive the same input image, and their respective encoders produce features $\mathbf{Z}_{s}$ and $\mathbf{Z}_{t}$ respectively. A learnable linear projection matrix $\mathbf{P}$ is used to project $\mathbf{Z}_{t}$ to the dimensions of $\mathbf{Z}_{s}$, giving $\bar{\mathbf{Z}}_{t}=\mathbf{Z}_{t}\mathbf{P}$. A distance function $d$ is then used between the student features and the projected teacher features:
@@
where $d$ can be any distance metric, such as the L2 loss used by FitNets [62] or the attention mapping described by AT [39]. In addition to this loss, we also use a task-specific supervision loss between the student model’s output $y_{s}$ and the ground truth labels $y$ to ensure the student’s output aligns with the target task. Since the teacher’s output is not used, we only perform a forward pass through its encoder in order to reduce the training compute required. The final loss is given by:
@@ 3.4 Decoupled Feature Distillation
$\leq\color[rgb]{0.265625,0.0078125,0.3359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.0078125,0.3359375}\pgfsys@color@rgb@stroke{0.265625}{0.0078125}{0.3359375}\pgfsys@color@rgb@fill{0.265625}{0.0078125}{0.3359375}\underbrace{\lVert\sum_{i\in k}\bar{\sigma}_{i}{\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{T}-\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}}\rVert_{2}}_{\text{knowledge transfer}}+\color[rgb]{0.1953125,0.7109375,0.47265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.7109375,0.47265625}\pgfsys@color@rgb@stroke{0.1953125}{0.7109375}{0.47265625}\pgfsys@color@rgb@fill{0.1953125}{0.7109375}{0.47265625}\underbrace{\lVert\sum_{i\notin k}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}\rVert_{2}}_{\text{regularisation}},$ Decoupled upper bound (5)
where $k$ denotes the set of indices indexing the non-zero singular values of $\bar{\mathbf{Z}}_{t}$. In practice, any metric $d$ that satisfies the triangle inequality has this decoupled upper bound in the cross-task setting. This result shows that the distillation loss can be decomposed into a knowledge transfer and an implicit spectral regularisation component. It explains how the inverted projection can help to improve performance even when there is little or no knowledge to transfer from the teacher: through a low-rank regularisation on the feature space.
@@ 3.5 Teacher-Free Distillation
The decoupled feature distillation in equation 5 allows us to introduce a novel spectral regularisation loss $\mathcal{L}_{spectral}$. This loss captures the regularisation effect of the cross-task distillation process without the use of any teacher, therefore we call this method "teacher-free distillation". It minimises the least-significant singular vectors of the student model’s features while keeping the most-significant. We define the spectral regularisation loss as follows. Assuming that the singular values/vectors are sorted from most to least significant, the loss can be given as follows:
@@ 4 Experiments and Results
In the following sections, we first give the implementation details and then present the results for the specific student tasks. In all of these setups, the teacher model is frozen, and the inverted projection and student model are jointly trained end-to-end using the loss function shown in equation 2.
Table 2: Comparison for different cross-task settings. We observe that our inverted projector is effective across many different task pairs and even for the same-task distillation settings.
(a) Semantic segmentation
@@ 4.1 Implementation details
The general framework used is described in section 3, and shown in figure 0(b). The model architectures used for the student and teacher vary depending on the task pairs. As an example, our depth estimation student is an encoder-decoder architecture with either a MobileNetV2, ResNet50, or EfficientNet-B0 backbone, and the frozen teacher model is a ViT-B/32 [18] trained for classification or the SwinV2-B [43] backbone of AiT [53] pretrained for instance segmentation. All architectures used follow an encoder-decoder structure. The student and teacher features, $\mathbf{Z}_{s}$ and $\mathbf{Z}_{t}$, are extracted immediately after the encoder of the model in question. All decoders used require features with a spatial (height and width) dimension, therefore if the teacher model’s encoder has a final pooling layer (as in the case of the classification teacher), this is removed. Full architectural details for the models used for each task pair are included in the supplementary material.
@@ 4.2 Monocular depth estimation
Monocular depth estimation is the task of inferring the depth, or distance to the camera, of every point shown in a single image. It is a challenging problem, as there is a many-to-one mapping from 3D scenes to a given 2D depth-image, but nonetheless deep learning methods have been able to perform very well by learning suitable priors over the data.
When experimenting with depth estimation as a target task, we make use of teachers trained on tasks that range from similar to dissimilar to the target task of depth estimation, in terms of the overlap in knowledge. We use a depth teacher for the same-task distillation, and the instance segmentation and classification tasks for the increasingly dissimilar teacher tasks. Finally, we use a randomly initialised and frozen teacher for the most extreme cross-task setting. Experiments are run on the NYUv2 dataset [66].
Our results are shown in table 3.2. As expected (see section 3), the use of our inverted projection produces improvements in performance when using dissimilar teacher tasks (random and classification), and gives similar or worse performance when the teacher’s task is similar to the target (instance segmentation and depth estimation teachers).
We use four different feature distillation methods from the same-task literature to show the utility of our method as a drop-in for use in the cross-task setting: FitNets [61], Attention Transfer (AT) [80], Probabilistic Knowledge Transfer (PKT) [54], and Ensemble (the projector ensemble method of [15]). The cross-task improvement using the inverted projection is most pronounced with FitNets and Ensemble, but some improvement can also be seen on AT and PKT. All teachers produce improvement over the baseline (where no teacher is used and the student is trained in isolation) with almost all methods. This indicates that in the case where it is uncertain the degree to which the teacher and student tasks are dissimilar, the use of our inverted projection will in general not be harmful.
[Input]                           [Ground Truth]                           [Depth Teacher]                           [Insseg Teacher]                           [Random Teacher]
[]Refer to caption []Refer to caption []Refer to caption []Refer to caption []Refer to caption []Refer to caption []Refer to caption []Refer to caption []Refer to caption []Refer to caption
Figure 2: Qualitative results on NYUv2 (depth) using different teacher tasks: results from depth estimation, instance segmentation, and randomly-initialised teachers on a MobileNetV2 [24] student. In each case, we use the optimal projection type for the teacher task: the depth teacher’s task-specific knowledge is desired, so we use a traditional projection, whereas the instance segmentation and random teachers both have irrelevant knowledge that must be discarded and so perform best with our novel inverted projection, which is able to remove the irrelevant features if needed.
We also provide qualitative examples using same-task, similar-task, and randomly-initialised teachers on a different depth estimation student, shown in figure 2. In all cases, we are able to obtain qualitatively good performance.
@@ 4.3 Semantic segmentation
Semantic segmentation is the task of labelling every pixel in the input image with a class. Our experiments are performed using MSCOCO [41], an 80-class segmentation dataset. We validate the effectiveness of our inverted projector using segmentation, classification, or randomly initialised teachers. In all experiments, we use a simple L2 loss between the projected teacher features and the student features. The results can be seen in table 2(a). We are able to obtain significant improvement with all the teacher tasks considered, with the best improvements seen with the random teacher. This follows: both classification and semantic segmentation have significant overlap in knowledge, but the random teacher has significant task-irrelevant information that our inverted projection is able to discard. This further empirically validates the regularisation components described in equation 5.
@@ 4.4 Image-to-image translation
We experiment with two image-to-image translation tasks: transforming satellite images into maps, and colourisation of black-and-white images. These two student tasks are particularly important, as they do not have significant knowledge overlap with any of the teacher tasks used: classification, depth estimation, instance segmentation, and the randomly-initialised teacher. In contrast, segmentation and classification both share a common goal of understanding the semantic context of the world, while depth estimation and segmentation have been shown to aid one another in the literature (see section 2). The results are shown in tables 2(c) and 2(b). Given the relative dissimilarity of the student tasks with all teacher tasks, it is not surprising that our inverted projection performs well in all cases. We are able to use our inverted projection to produce a significant improvement over the baseline with all teacher tasks, including with the randomly initialised teacher.
(a) $\blacksquare$ Seg. $\rightarrow$ Seg.
(b) $\blacksquare$ Class. $\rightarrow$ Seg.
(c) $\blacksquare$ Random $\rightarrow$ Seg.
Figure 3: Evolution of singular values of the projection matrix $\mathbf{P}$ under different cross-task settings and projector types. An L2 loss is used. Green area highlights the rank of $\mathbf{P}$. The projection tends towards a higher rank either when using the traditional projection or when using a same-task or similar-task teacher. The low-rank when using our inverted projection in the cross-task setting allows irrelevant features to be filtered out, if necessary for the task pair. Top row: traditional projection, Bottom row: our inverted projection. Numerical rank is used with a tolerance set to $\sigma_{1}*0.01$.
4.5 Training Dynamics of the Inverted Projector
By observing the singular value spectrum of the projector weights and how they evolve over the course of training, we are able to provide further insight into the role of our novel inverted projection for cross-task distillation, as compared to the traditional projection. Figure 3.2 shows the singular value spectrum of the projector weights throughout the training process of a segmentation student with both similar and different teacher tasks, using either the traditional projection or our novel inverted projection. It can be seen that the traditional projection fails to disregard many of the less-dominant singular values, leading to a higher-rank projection in general. As discussed in section 3.2, this is especially detrimental when there is a significant task gap, as it encourages the student to learn task-irrelevant features. When the student model is small, this can significantly degrade the target task performance. However, when using our inverted projection, we observe a consistently lower rank across training for all tasks, compared to the traditional projection. This is because of the inverted projection’s ability to suppress the task-irrelevant singular vectors from the teacher model: while the traditional projection remains consistently high-rank regardless of the dissimilarity of the student and teacher tasks, our inverted projection is able to adapt to discard the increasing quantity of undesirable task-specific knowledge encoded in the increasingly dissimilar teacher features.
4.6 Ablation Study
(a) Depth $\rightarrow$ Depth
(b) Insseg $\rightarrow$ Depth
(c) Random $\rightarrow$ Depth
Figure 4: Comparing performance of different-sized depth students with both the traditional projection and our novel inverted projection. Where there is knowledge to transfer from teacher to student (i.e. the two tasks are similar), the traditional projection performs better, but when the teacher is random, the opposite is true. Only decoder size is varied. A MobileNetV2 [24] is used as the backbone.
@@ 4.6 Ablation Study
Teacher-free distillation. As shown in sections 4.2, 4.3, and 4.4, we are able to obtain significant performance improvements even when the teacher is randomly initialised and then frozen, thus containing no task-specific knowledge at all. This reinforces the conclusion reached in section 3.4: the distillation loss function $\mathcal{L}_{distill}$ may be comprised of a knowledge transfer component and a spectral regularisation component. In the case where there is no knowledge to transfer between the teacher and the student, only a regularising effect can explain the performance improvement over the baseline.
Table 3: Teacher-free distillation using our spectral regularisation loss on the NYUv2 dataset using AiT [53] on a SwinV2-b base. The regularisation loss is generally robust to different values of $r$, with $r=2$ being optimal.
Method RMS $\downarrow$ Abs $\downarrow$ $\delta_{1}$uparrow
(Baseline) AiT (SwinV2-B) [53] 0.365 0.105 0.907
$\mathcal{L}_{spectral}$ $(r=1)$ 0.352 0.105 0.902
$\mathcal{L}_{spectral}$ $(r=2)$ 0.340 0.096 0.914
$\mathcal{L}_{spectral}$ $(r=4)$ 0.349 0.099 0.911
$\mathcal{L}_{spectral}$ $(r=8)$ 0.344 0.096 0.910
$\mathcal{L}_{spectral}$ $(r=16)$ 0.348 0.098 0.912
$\mathcal{L}_{spectral}$ $(r=32)$ 0.347 0.100 0.909
To control for this, and to provide further evidence of the loss decoupling described in equation 5, we perform experiments using the spectral regularisation loss defined in equation 6. We experiment with different values of $r$ to vary the strength of the loss using a depth estimation student model to be trained on NYUv2 [66] with no teacher model. The results are shown in table 3, and show that our spectral regularisation loss improves performance significantly over the baseline for all values of $r$ tried, with $r=2$ producing the greatest improvement. This finding is significant, as it gives additional support to the decoupling of $\mathcal{L}_{distill}$ into knowledge transfer and regularisation terms (as shown in equation 5), and further validates the use of the randomly-initialised teacher in our comparisons. Additionally, it provides a useful drop-in tool that can enhance performance with little or no additional cost. Since our proposed regularisation loss is an upper bound on feature distillation, an interesting thought is how it compares to other distillation methods. In table 4 we consider knowledge distillation on the large-scale ImageNet-1K dataset, and we observe that our simple regularisation loss achieves competitive performance with many state-of-the-art knowledge distillation methods.
Table 4: Comparing our novel teacher-free spectral regularisation loss to other state-of-the-art KD methods on ImageNet-1K [16]. Top row is the teacher model used by the KD methods that use a teacher. All methods use a DeiT-Ti [71] student model, with DeiT-Ti[Uncaptioned image] describing the distilled variant using distillation tokens.
@@
RegNety 160 [57] 82.6 84M
@@
DeiT-Ti[Uncaptioned image] [71] 74.5 6M
Co-advise [60] 74.9 6M
DearKD  [11] 74.8 6M
USKD [77] 75.0 6M
@@
DeiT-Ti [71] 72.2 5M
@@
Different architecture pairs. To demonstrate the generality of our proposed inverted projection in various cross-task settings, we perform an ablation across several differently-sized student models with similar and dissimilar task pairs. Figure 4 shows that the performance drop or improvement is consistent for both the very small and moderately large student models, across different task pairs. It also mirrors the findings of our previous experiments in sections 4.2, 4.3, and 4.4, showing that the similarity of the teacher and student tasks matters, and that our novel inverted projection performs best when the two tasks are dissimilar.
Table 5: Comparisons with different architecture pairs. All experiments perform cross-task distillation from a classification teacher to a depth estimation student. The inverted projection is effective across various student model sizes and with different KD methods in this cross-task setting.
Student Arch (backbone) $\longrightarrow$ EfficientNet-B0 (5.3M) ResNet-50 (25.6M)
KD Method
Projection type
$\delta_{1}\uparrow$ Abs. $\downarrow$ RMS $\downarrow$ $\delta_{1}\uparrow$ Abs. $\downarrow$ RMS $\downarrow$
None (baseline) N/A 0.845 0.127 0.440 0.811 0.144 0.480
AT [80]
Original
0.850 0.125 0.433 0.814 0.143 0.477
Inverted (ours)
0.853 0.123 0.430 0.816 0.143 0.475
Improvement 0.35% 1.61% 0.79% 0.26% 0.14% 0.38%
PKT [54]
Original
0.851 0.124 0.432 0.816 0.143 0.473
Inverted (ours)
0.853 0.123 0.431 0.821 0.141 0.470
Improvement 0.25% 1.29% 0.30% 0.63% 1.19% 0.70%
FitNets [62]
Original
0.845 0.125 0.439 0.812 0.145 0.480
Inverted (ours)
0.850 0.124 0.434 0.813 0.142 0.476
Improvement 0.50% 0.53% 1.34% 0.16% 1.93% 0.77%
Table 5 shows results distilling from a classification teacher to a depth student using two student backbones of significantly-different sizes: an EfficientNet-B0 (5.3M params) and a ResNet-50 (25.6M params). These students are chosen to illustrate both a small and a large capacity gap between the student and teacher models. Three different KD methods are used. Our inverted projector outperforms the traditional projector across all metrics for all three KD methods and both student backbone architectures in this cross-task setting, thus demonstrating the generality of our inverted projector for a variety of practical KD settings.

### v2

@@ 3.3 Setup and Training loss
Our cross-task knowledge distillation pipeline is shown in figure LABEL:fig:overview-ours-cross-task. It consists of a trainable student model, which is to be trained on a given target task, and a frozen teacher model, which is pre-trained on a different task. This setup is in contrast to the traditional same-task knowledge distillation setting, which is shown in figure LABEL:fig:overview-same-task. Both the student and teacher models receive the same input image, and their respective encoders produce features $\mathbf{Z}_{s}$ and $\mathbf{Z}_{t}$ respectively. A learnable linear projection matrix $\mathbf{P}$ is used to project $\mathbf{Z}_{t}$ to the dimensions of $\mathbf{Z}_{s}$, giving $\bar{\mathbf{Z}}_{t}=\mathbf{Z}_{t}\mathbf{P}$. A distance function $d$ is then used between the student features and the projected teacher features:
@@
where $d$ can be any distance metric, such as the L2 loss used by FitNets [Romero et al.(2015b)Romero, Ballas, Kahou, Chassang, Gatta, and Bengio] or the attention mapping described by AT [Le et al.(2020)Le, Vo, and Thoa]. In addition to this loss, we also use a task-specific supervision loss between the student model’s output $y_{s}$ and the ground truth labels $y$ to ensure the student’s output aligns with the target task. Since the teacher’s output is not used, we only perform a forward pass through its encoder in order to reduce the training compute required. The final loss is given by:
@@ 3.4 Decoupled Feature Distillation
$\leq\color[rgb]{0.265625,0.0078125,0.3359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.265625,0.0078125,0.3359375}\underbrace{\lVert\sum_{i\in k}\bar{\sigma}_{i}{\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{T}-\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}}\rVert_{2}}_{\text{knowledge transfer}}+\color[rgb]{0.1953125,0.7109375,0.47265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.7109375,0.47265625}\underbrace{\lVert\sum_{i\notin k}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{T}\rVert_{2}}_{\text{regularisation}},$ Decoupled upper bound (5)
where $k$ denotes the set of indices indexing the non-zero singular values of $\bar{\mathbf{Z}}_{t}$. In practice, any metric $d$ that satisfies the triangle inequality has this decoupled upper bound in the cross-task setting. This result shows that the distillation loss can be decomposed into a knowledge transfer and an implicit spectral regularisation component. It explains how the inverted projection can help to improve performance even when there is little or no knowledge to transfer from the teacher: through a low-rank regularisation on the feature space. We empirically observe this emergent decoupling in figure LABEL:fig:eigplots. Here we see that an inverted projected is more effective at removing irrelevant task information, which is important in the cross-task KD setting.
@@ 3.5 Teacher-Free Distillation
The decoupled feature distillation in equation 5 allows us to introduce a novel spectral regularisation loss $\mathcal{L}_{spectral}$. This loss captures the regularisation effect of the cross-task distillation process without the use of any teacher, therefore we call this method "teacher-free distillation" as is similarly done in other works [Yuan et al.(2020)Yuan, Tay, Li, Wang, and Feng]. Its objective is to minimise the least-significant singular vectors of the student model’s features while keeping the most-significant. We define the spectral regularisation loss as follows. Assuming that the singular values/vectors are sorted from most to least significant, the loss can be given as follows:
@@ 4 Experiments and Results
Satellite-to-map conversion
Table 2: Comparison for different cross-task settings. We observe that our inverted projector is effective across many different task pairs and even for the same-task settings.
@@ 4.1 Implementation details
The general framework used is described in section 3, and shown in figure LABEL:fig:overview-ours-cross-task. The model architectures used for the student and teacher vary depending on the task pairs. As an example, our depth estimation student is an encoder-decoder architecture with either a MobileNetV2, ResNet50, or EfficientNet-B0 backbone, and the frozen teacher model is a ViT-B/32 [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] trained for classification or the SwinV2-B [Liu et al.(2022)Liu, Hu, Lin, Yao, Xie, Wei, Ning, Cao, Zhang, Dong, Wei, and Guo] backbone of AiT [Ning et al.(2023)Ning, Li, Zhang, Geng, Dai, He, and Hu] pretrained for instance segmentation. All architectures used follow an encoder-decoder structure. The student and teacher features, $\mathbf{Z}_{s}$ and $\mathbf{Z}_{t}$, are extracted immediately after the encoder of the model in question. All decoders used require features with a spatial (height and width) dimension, therefore if the teacher model’s encoder has a final pooling layer (as in the case of the classification teacher), this is removed.
@@ 4.2 Monocular depth estimation
Monocular depth estimation is the task of inferring the depth, or distance to the camera, of every point shown in a single image. It is a challenging problem, as there is a many-to-one mapping from 3D scenes to a given 2D depth-image. When experimenting with depth estimation as a target task, we make use of teachers trained on tasks that range from similar to dissimilar to the target task of depth estimation, in terms of the overlap in knowledge. We use a depth teacher for the same-task distillation, and the instance segmentation and classification tasks for the increasingly dissimilar teacher tasks. Finally, we use a randomly initialised and frozen teacher for the most extreme cross-task setting. Experiments are run on the NYUv2 dataset [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus]. Our results are shown in table 3.2. As expected (see section 3), the use of our inverted projection produces improvements in performance when using dissimilar teacher tasks (random and classification), and gives similar or worse performance when the teacher’s task is similar to the target (instance segmentation and depth estimation teachers).
We use four different feature distillation methods from the same-task literature to show the utility of our method as a drop-in for use in the cross-task setting: FitNets [Romero et al.(2015a)Romero, Ballas, Ebrahimi Kahou, Chassang, Gatta, and Bengio], Attention Transfer (AT) [Zagoruyko and Komodakis(2019)], Probabilistic Knowledge Transfer (PKT) [Passalis and Tefas(2018)], and Ensemble (the projector ensemble method of [Chen et al.(2022b)Chen, Wang, Liu, Xu, de Hoog, and Huang]). The cross-task improvement using the inverted projection is most pronounced with FitNets and Ensemble, but there is some improvement with AT and PKT.
@@ 4.3 Semantic segmentation
Semantic segmentation is the task of labelling every pixel in the input image with a class. Our experiments are performed using MSCOCO [Lin et al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick], an 80-class segmentation dataset. We validate the effectiveness of our inverted projector using segmentation, classification, or randomly initialised teachers. In all experiments, we use a simple L2 loss between the projected teacher features and the student features. Results shown in table 2. We are able to obtain significant improvement with all the teacher tasks considered, with the best improvements seen with the random teacher. This follows: both classification and semantic segmentation have significant overlap in knowledge, but the random teacher has significant task-irrelevant information that our inverted projection is able to discard. This further empirically validates the regularisation components described in equation 5.
@@ 4.4 Image-to-image translation
We experiment with two image-to-image translation tasks: transforming satellite images into maps, and colourisation of black-and-white images. These two student tasks are particularly important, as they do not have significant knowledge overlap with any of the teacher tasks used: classification, depth estimation, instance segmentation, and the randomly-initialised teacher. In contrast, segmentation and classification share a common goal of understanding semantic context of the world, while depth estimation and segmentation have been shown to aid one another (see section 2). Results are shown in table 2. Given the relative dissimilarity of the student tasks with all teacher tasks, unsurprisingly, our inverted projection performs well in all cases. We are able to use our inverted projection to produce a significant improvement over the baseline with all teacher tasks, including with the randomly initialised teacher.
4.5 Teacher-free distillation.
@@
As shown in sections 4.2, 4.3, and 4.4, we are able to obtain significant performance improvements even when the teacher is randomly initialised and then frozen, thus containing no task-specific knowledge at all. This reinforces the conclusion reached in section 3.4: the distillation loss function $\mathcal{L}_{distill}$ may be comprised of a knowledge transfer component and a spectral regularisation component. In the case where there is no knowledge to transfer between the teacher and the student, only a regularising effect can explain the performance improvement over the baseline. To control for this, and to provide further evidence of the loss decoupling described in equation 5, we perform experiments using the spectral regularisation loss (equation 6). Experimenting with different $r$ values in a depth estimation model trained on NYUv2 [Silberman et al.(2012)Silberman, Hoiem, Kohli, and Fergus] without a teacher, we find that spectral regularisation significantly enhances performance across all $r$ values, particularly at $r=2$ (see appendix). This supports the decoupling of $\mathcal{L}_{distill}$ into knowledge transfer and regularisation terms (equation 5) and further validates using a randomly-initialised teacher. In table 3 we consider knowledge distillation on the large-scale ImageNet-1K dataset, and we observe that our simple regularisation loss achieves competitive performance with many state-of-the-art knowledge distillation methods.
@@
RegNety 160 [Radosavovic et al.(2020)Radosavovic, Kosaraju, Girshick, He, and Dollár] 82.6 84M
@@
DeiT-Ti[Uncaptioned image] [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 74.5 6M
Co-advise [Ren et al.(2022)Ren, Gao, Hua, Xue, Tian, He, and Zhao] 74.9 6M
DearKD  [Chen et al.(2022a)Chen, Cao, Zhong, Zhang, Gao, and Tao] 74.8 6M
USKD [Yang et al.(2023)Yang, Zeng, Li, Zhang, Yuan, and Li] 75.0 6M
@@
DeiT-Ti [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 72.2 5M
@@
Table 3: Comparing our novel teacher-free spectral regularisation loss to other state-of-the-art KD methods on ImageNet-1K [Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei]. Top row is the teacher model used by the KD methods that use a teacher. All methods use a DeiT-Ti [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] student model, with DeiT-Ti[Uncaptioned image] describing the distilled variant using distillation tokens.